OpenStudy (anonymous):

I NEED HELP PLEASE Choose the correct rewriting of this algebraic fraction and give the exclusion 4Y(8Y+24)^-1

OpenStudy (anonymous):

A;Y/2(2)(Y+3) B;2Y/2Y+3

OpenStudy (anonymous):

I THINK A

OpenStudy (anonymous):

4y/(8y+24) = y/(2y+6)=y/(2(y+3)) y can not equal -3

OpenStudy (anonymous):

THANKS